OpenStudy (anonymous):

Gaskin and sloan raced tpo the cotton patch. Gaskins speed was 12 kilometers per hour, while sloans was only 8 kilometers per hour. What was the time of each if gaskins time was 5 hours less than sloans time?

OpenStudy (barrycarter):

Let x = distance to cotton patch. Then solve: x/12 = x/8 -5

OpenStudy (amistre64):

Gaskin and sloan raced tpo the cotton patch. so this is linear in nature Gaskins speed was 12 kilometers per hour, while sloans was only 8 kilometers per hour. What was the time of each if gaskins time was 5 hours less than sloans time?

OpenStudy (anonymous):

somebody give me an equation

OpenStudy (amistre64):



OpenStudy (amistre64):

an equation is something we have to build from the information given ....

OpenStudy (anonymous):

ik

OpenStudy (amistre64):

so, distance remains the same in each case since the cotton field hasnt moved

OpenStudy (amistre64):

d = speed.g*time.g d = speed.s*time.s are how we should relate this things

OpenStudy (amistre64):

speed.g = 12 ; speed.s=8 d = 12*time.g d = 8*time.s time.g = time.s-5 d = 12*(time.s-5) d = 8*time.s since d=d we equate the 2 to find the time for sloan

OpenStudy (anonymous):

ummm could u put it in an equation? like the full one

OpenStudy (amistre64):

there are a myriad of different ways to solve this; im not sure which "equation" you are suggesting.

OpenStudy (anonymous):

just one that gets the answers

OpenStudy (amistre64):

well, if we equate the ds .... since d=d d = d 12(s-5) = 8(s)

OpenStudy (anonymous):

good is that the whole thing?

OpenStudy (amistre64):

how would I know? I have to take a blind stab at this thing just to try to make sure im even answering the right question; much less try to determine what level of math you are trying to solve it in ..... but as is, you should be able to work out the time for sloan with that and compare it to gs time

OpenStudy (anonymous):

its algebra 1

OpenStudy (amistre64):

then yes, that should be the "whole thing" :)

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

solve it the way an algebra student would

OpenStudy (anonymous):

can u solve it?

OpenStudy (anonymous):

im done dude u can get off now